# ASSESS: A Semantic and Structural Evaluation Framework for Statement Similarity

### v1

@@ Appendix A Proof of Theorem 1
$0\in\mathcal{F}$ so $\mathcal{F}\neq\emptyset$. Let $\widehat{f}:X\times X\to[0,\infty]$ be a function defined as
@@
It suffices to show that $\widehat{f}\in\mathcal{F}$. By definition of $\widehat{f}$, for any point pairs $(x,y)\in X\times X$ and any neighbourhood $\mathcal{N}$ of $\widehat{f}(x,y)$, there exists $f_{x,y,\mathcal{N}}\in\mathcal{F}$ such that
@@
Next, we show that $\widehat{f}$ satisfies the extra requirements in the definition of $\mathcal{F}$. Since for any $f\in\mathcal{F}$, $f(x,y)\leq b(x,y)$, we have
$\widehat{f}(x,y)\leq b(x,y).$
Take any $((x,y),(u,v))\in T$. Notice that for any $f\in\mathcal{F}$, $f(x,y)\leq f(u,v)$. Now for any neighbourhood $\mathcal{N}$ of $\widehat{f}(x,y)$, we have
@@
Appendix B Pseudocode for TransTED
This section provides the detailed pseudocode for our TransTED algorithm. The implementation consists of two main functions: a primary wrapper TransTED that handles initial checks, and the core TED AFTER TRANSFORMATION function, which executes the heuristic search. The search (line 13) is guided by using the Tree Edit Distance as its heuristic (line 18). If a proof of equivalence is found, the distance is 0; otherwise, if the time limit is exceeded, the algorithm returns the smallest TED found among all visited nodes (line 21).
@@ Appendix B Pseudocode for TransTED
10:  eq $\leftarrow$ UseTactic(eq, norm_num1)
11:  if Completed(eq) $\lor$ Completed(UseTactic(eq, rfl)) then
12:   return 0
13:  end if
14:  HeuristicSearch(
15:   start : eq,
16:   stop : the goal completed $\lor$ time limit exceeded (TLE)
17:   search method : suggestions by rw? and some given additional tactic commands
18:   valid nodes: a single goal of equality, or “completed”
19:   heuristic function : TED between the expressions trees of both sides of the equality
20:  )
21:  if TLE then
22:   return smallest TED of all the visited nodes
23:  else
24:   return 0 $\triangleright$ The goal is completed within the time limit
25:  end if
26:end function
Appendix C Experimental Details and Case Studies
C.1 Detailed Experimental Results
This section provides a detailed breakdown of the experimental results on the EPLA-miniF2F and EPLA-ProofNet datasets in Table 4 and Table 5, respectively. The tables report the number of True Positives (TP), True Negatives (TN), False Positives (FP), and False Negatives (FN) for each evaluated metric.
@@ C.1 Detailed Experimental Results
TP TN FP FN Precision Recall Accuracy Kappa
Identity Match 27 97 0 249 100.00% 9.78% 33.24% 0.05
Typecheck 276 0 97 0 73.99% 100.00% 73.99% 0.00
BLEU 174 77 20 102 89.69% 63.04% 67.29% 0.33
Majority Voting 170 78 19 106 89.95% 61.59% 66.49% 0.33
Definitional Equality 92 96 1 184 98.92% 33.33% 50.40% 0.20
BEq 135 97 0 141 100.00% 48.91% 62.20% 0.33
TED Similarity 206 63 34 70 85.83% 74.64% 72.12% 0.35
TransTED Similarity 235 59 38 41 86.08% 85.14% 78.82% 0.46
@@
TP TN FP FN Precision Recall Accuracy Kappa
Identity Match 1 71 0 79 100.00% 1.25% 47.68% 0.01
Typecheck 80 0 71 0 52.98% 100.00% 52.98% 0.00
BLEU 61 37 34 19 64.21% 76.25% 64.90% 0.29
Majority Voting 51 52 19 29 72.86% 63.75% 68.21% 0.37
Definitional Equality 8 69 2 72 80.00% 10.00% 50.99% 0.07
BEq 23 71 0 57 100.00% 28.75% 62.25% 0.28
TED Similarity 64 38 33 16 65.98% 80.00% 67.55% 0.34
TransTED Similarity 70 37 34 10 67.31% 87.50% 70.86% 0.40
C.2 Illustrative Examples of TransTED Similarity
@@
Proof given by TransTED:
@@
# mathd_algebra_142 (from miniF2F)
@@
Proof given by TransTED:
@@
C.3 False Positives in Definitional Equality
This case study investigates the root cause of the false positives recorded by the Definitional Equality metric—notably, one on the EPLA-miniF2F and two on EPLA-ProofNet.
# mathd_algebra_176
@@ C.3 False Positives in Definitional Equality
theorem thm_P (x : $\mathbb{R}$) : (x + 1) ^ 2 * x = x ^ 3 + 2 * x ^ 2 + x := by sorry
@@
theorem thm_Q {R : Type*} [CommRing R] (x : R) : (x + 1) ^ 2 * x = x ^ 3 + 2 * x ^ 2 + x := by sorry
@@
# exercise_1_3_8
Label:
theorem thm_P : Infinite (Equiv.Perm $\mathbb{N}$) := by sorry
Prediction:
theorem thm_Q $\mathrm{\Omega}$ : Type u_1} [Infinite $\mathrm{\Omega}$] : Infinite (Equiv.Perm $\mathrm{\Omega}$) := by sorry
Definitional Equality
example : thm_P = thm_Q := by rfl
# exercise_9_4_2c
NL: Prove that $x^{4}+4x^{3}+6x^{2}+2x+1$ is irreducible in $\mathbb{Z}[x]$.
Label:
theorem thm_P : Irreducible (X^4 + 4*X^3 + 6*X^2 + 2*X + 1 : Polynomial $\mathbb{Z}$) := by sorry
Prediction:
theorem thm_Q : Irreducible (wilsons_poly : $\mathbb{Z}$[X]) := by sorry
Definitional Equality
example : thm_P = thm_Q := by rfl
Appendix D Prompt Templates
This section presents the prompts for the majority voting phase, which uses InternLM2-Math-Plus-7B (Ying et al., 2024b) for back-translation and DeepSeek-V3.1 (Liu et al., 2024) for NLI Check.
[…]

### v2

@@ Appendix C Proof of Theorem 1
$0\in\mathcal{F}$ so $\mathcal{F}\neq\emptyset$. Let $\widehat{f}:X\times X\to\mathbb{R}_{\geq 0}$ be a function defined as
@@
Since $f(x,y)\leq b(x,y)$ for any $f\in\mathcal{F}$, $\widehat{f}(x,y)\leq b(x,y)<\infty$. It suffices to show that $\widehat{f}\in\mathcal{F}$. By definition of $\widehat{f}$, for any point pairs $(x,y)\in X\times X$ and any neighbourhood $\mathcal{N}$ of $\widehat{f}(x,y)$, there exists $f_{x,y,\mathcal{N}}\in\mathcal{F}$ such that
@@
Next, we show that $\widehat{f}$ satisfies the extra requirements in the definition of $\mathcal{F}$. We have shown that $\widehat{f}(x,y)\leq b(x,y)$. Now take any $((x,y),(u,v))\in T$. Notice that for any $f\in\mathcal{F}$, $f(x,y)\leq f(u,v)$. Now for any neighbourhood $\mathcal{N}$ of $\widehat{f}(x,y)$, we have
@@
Appendix D Pseudocode for TransTED
This section provides the detailed pseudocode for our TransTED algorithm. The implementation consists of two primary functions: a primary wrapper TransTED that handles initial checks, and the core TEDAfterTransformation function, which executes the heuristic search. The search (line 13) is guided by using the Tree Edit Distance as its heuristic (line 18). If a proof of equivalence is found, the distance is 0; otherwise, if the time limit is exceeded, the algorithm returns the smallest TED found among all visited nodes (line 21).
@@ Appendix D Pseudocode for TransTED
10:  if Completed(UseTactic(eq, exact?)) then
11:   return 0
12:  end if
13:  HeuristicSearch(
14:   start : eq,
15:   stop : the goal completed $\lor$ node limit exceeded (NLE) $\lor$ time limit exceeded (TLE)
16:   search method : suggestions by rw? and some given additional tactic commands
17:   valid nodes: a single goal of equality, or “completed”
18:   heuristic function : TED between the expressions trees of both sides of the equality
19:  )
20:  if TLE then
21:   return smallest TED of all the visited nodes
22:  else
23:   return 0 $\triangleright$ The goal is completed within the time limit
24:  end if
25:end function
Appendix E Benchmark Structure
The EPLA data format is structured as a tuple: (natural language statement, ground-truth formalization, candidate formalization, annotation label). The annotation label is further decomposed into three Boolean sub-labels, each designed to quantify a specific dimension of similarity and address a distinct application scenario.
Sub-label Definition Application Scenario
Provability Whether the two statements are semantically equivalent. To evaluate if a metric correctly assesses the semantic equivalence of two statements.
Likeness before transformation Whether the two (original) statements are structurally similar. To evaluate if a metric reflects the editing cost required to make one statement structurally identical to another.
Likeness after transformation Whether the two statements are structurally similar after semantic-preserving transformations To evaluate if a metric reflects the editing cost required to make one statement semantically identical to another.
Due to inherent logical dependencies between sub-labels, the label space is constrained to the five valid combinations. The underlying logic is governed by two key implications:
  • Provability $\implies$ Likeness after transformation: Since provable statements are inherently semantically equivalent, a valid transformation path exists by definition, satisfying the likeness criterion.
  • Likeness before transformation $\implies$ Likeness after transformation: Structural similarity serves as a strict condition; the editing cost required to demonstrate semantic equivalence is naturally upper-bounded by the cost required to achieve structural identity.
Provability Likeness before transformation Likeness after transformation
True True True
True False True
False True True
False False True
False False False
Appendix F Experimental Details and Case Studies
This appendix supplements the main paper with comprehensive experimental details and qualitative case studies. We begin by outlining the methodology for threshold selection in Appendix F.1, followed by the detailed experimental results in Appendix F.2. We then provide a qualitative assessment of metric performance, presenting illustrative examples of TransTED Similarity in Appendix F.3 and discussing false positives in Definitional Equality and BEq in Appendix F.4. Finally, the specific prompt templates used for autoformalization and evaluation are documented in Appendix F.5.
F.1 Threshold Selection
We observe that the choice of threshold is inherently benchmark-dependent. This dependency stems from the richness of the underlying mathematical formalism: plentiful mathematical structures $\Rightarrow$ more structurally variant expressions of the same meaning $\Rightarrow$ lower possible structural likeness.
For instance, the optimal TransTED Similarity threshold for EPLA-miniF2F ($0.61$) is significantly higher than that for EPLA-ProofNet ($0.38$). This disparity reflects the greater structural complexity and variation inherent to the latter. Consequently, we recommend that practitioners calibrate thresholds according to the structural diversity of their specific domain, using the optimal values reported in this work as reference baselines. To ensure robustness, we further advise evaluating model performance across a spectrum of thresholds rather than relying on a single static cut-off.
F.2 Detailed Experimental Results
Tables 6 and 7 present the experimental results on the EPLA-miniF2F and EPLA-ProofNet benchmarks, respectively. For each metric, we report a detailed classification breakdown (TP, TN, FP, FN) alongside computational costs, measured by total and average execution times. Additionally, Figures 5 and 6 analyze the tactic command counts for TransTED Similarity.
@@ F.2 Detailed Experimental Results
TP TN FP FN Precision Recall Accuracy Kappa Tot. Time (s) Avg. Time (s)
Identity Match 54 217 0 560 100.00% 8.79% 32.61% 0.05 0.1240 0.0001
BLEU 454 119 98 160 82.25% 73.94% 68.95% 0.26 10.6390 0.0128
Majority Voting 192 198 19 422 91.00% 31.27% 46.93% 0.14 11343.1519 13.6500
Definitional Equality 200 211 6 414 97.09% 32.57% 49.46% 0.19 215.8885 0.2598
BEq 281 213 4 333 98.60% 45.77% 59.45% 0.29 26133.5967 31.4484
TED Similarity 442 136 81 172 84.51% 71.99% 69.55% 0.31 998.8283 1.2020
TransTED Similarity 424 159 58 190 87.97% 69.06% 70.16% 0.35 42028.5440 50.5759
@@
TP TN FP FN Precision Recall Accuracy Kappa Tot. Time (s) Avg. Time (s)
Identity Match 9 172 0 235 100.00% 3.69% 43.51% 0.03 0.1179 0.0003
BLEU 107 131 41 137 72.30% 43.85% 57.21% 0.18 5.7977 0.0139
Majority Voting 110 117 55 134 66.67% 45.08% 54.57% 0.12 5371.1336 12.9114
Definitional Equality 25 168 4 219 86.21% 10.25% 46.39% 0.07 114.6867 0.2757
BEq 80 171 1 164 98.77% 32.79% 60.34% 0.28 14872.7313 35.7518
TED Similarity 199 70 102 45 66.11% 81.56% 64.67% 0.23 549.1361 1.3200
TransTED Similarity 200 80 92 44 68.49% 81.97% 67.31% 0.30 54343.4060 130.6332
Figure 5: Statistic of the tactic commands in EPLA-miniF2F
Figure 6: Statistic of the tactic commands in EPLA-ProofNet
F.3 Illustrative Examples of TransTED Similarity
@@
Proof given by TransTED Similarity:
@@
EPLA-miniF2F #170
@@
Proof given by TransTED Similarity:
@@
F.4 False Positives in Definitional Equality and BEq
This case study investigates the root cause of the false positives (FP) recorded by the Definitional Equality metric and the BEq metric, classified into three categories according to causes.
Causes miniF2F ProofNet Definitional Equality FP BEq FP
Implicit Variable Matching #24, #701 #115, #128, #402 ++ -
Over-simplifying #378, #603, #639, #641 - ++ ++
Header Mixing - #280 ++ ++
F.4.1 Implicit Variable Matching
EPLA-miniF2F #24
@@ F.4.1 Implicit Variable Matching
theorem mathd_algebra_176 (x : $\mathbb{R}$) : (x + 1) ^ 2 * x = x ^ 3 + 2 * x ^ 2 + x := by sorry
@@
theorem my_favorite_theorem {R : Type*} [CommRing R] (x : R) : (x + 1) ^ 2 * x = x ^ 3 + 2 * x ^ 2 + x := by sorry
Definitional Equality
example : mathd_algebra_176 = my_favorite_theorem := by rfl
EPLA-miniF2F #701
NL: Expand (x+3)(2x6)(x+3)(2x-6). Show that it is 2xˆ2-18.
Label:
theorem mathd_algebra_432 (x : $\mathbb{R}$) : (x + 3) * (2 * x - 6) = 2 * x ^ 2 - 18 := by sorry
Prediction:
theorem test_problem {R : Type*} [CommRing R] (x : R) : (x + 3) * (2 * x - 6) = 2 * x^2 - 18 := by sorry
Definitional Equality
example : mathd_algebra_432 = test_problem := by rfl
The error type here is identical to that in EPLA-miniF2F #24.
EPLA-ProofNet #115
Label:
theorem exercise_1_3_8 : Infinite (Equiv.Perm $\mathbb{N}$) := by sorry
Prediction:
theorem infinite_of_infinite_card_aux $\mathrm{\Omega}$ : Type u_1} [Infinite $\mathrm{\Omega}$] : Infinite (Equiv.Perm $\mathrm{\Omega}$) := by sorry
Definitional Equality
example : exercise_1_3_8 = infinite_of_infinite_card_aux := by rfl
EPLA-ProofNet #128
NL: Prove that $x^{4}+4x^{3}+6x^{2}+2x+1$ is irreducible in $\mathbb{Z}[x]$.
Label:
[…]
Prediction:
theorem irreducible_wilsons_poly : Irreducible (wilsons_poly : $\mathbb{Z}$[X]) := by sorry
Definitional Equality
example : exercise_9_4_2c = irreducible_wilsons_poly := by rfl
EPLA-ProofNet #402
NL: Let $f\colon X\rightarrow X$ be continuous. Show that if $X=[0,1]$, there is a point $x$ such that $f(x)=x$. (The point $x$ is called a fixed point of $f$.)
Label:
theorem exercise_24_3a [TopologicalSpace I] [CompactSpace I] (f : I $\rightarrow$ I) (hf : Continuous f) : $\exists$ (x : I), f x = x := by sorry
Prediction:
theorem test_problem (f : Set.Icc (0 : $\mathbb{R}$) 1 $\rightarrow$ Set.Icc (0 : $\mathbb{R}$) 1) (hf : Continuous f) : $\exists$ x, f x = x := by sorry
Definitional Equality
example : exercise_24_3a = test_problem := by rfl
In the Prediction, the model explicitly uses the concrete type, the interval $[0,1]$, for both the domain and codomain of the function $f$. However, in the Label, the theorem is stated more abstractly for any type $I$ equipped with needed instances. While $[0,1]$ can be endowed with such structure (making it a valid candidate for $I$), the type $0,1]\to[0,1]$ is not definitionally equal to the function type $I\to I$ for an abstract $I$. Therefore, the definitions are not considered identical by the kernel, the rfl tactic fails, and this case is correctly identified as a negative example.
F.4.2 Over-simplifying
EPLA-miniF2F #378
NL: What is the units digit of the sum of the squares of the first nine positive integers? Show that it is 5
Label:
theorem mathd_numbertheory_3 : ($\mathrm{\Sigma}$ x in Finset.range 10, (x + 1) ^ 2) % 10 = 5 := by sorry
Prediction:
theorem my_favorite_theorem : ($\mathrm{\Sigma}$ i in Finset.range 9, (i + 1)^2) % 10 = 5 := by sorry
These two theorems produce exactly the same numerical result. However, the natural language description specifies computing the squares of the first nine positive integers, while the Label computes the squares of the first ten positive integers. Although this does not change the final value, the semantics are different. Nevertheless, BEq reports that the two theorems are equivalent, which does not match the actual situation.
EPLA-miniF2F #603
NL: What is the units digit of the sum of the squares of the first nine positive integers? Show that it is 5
Label:
theorem mathd_numbertheory_3 : ($\mathrm{\Sigma}$ x in Finset.range 10, (x + 1) ^ 2) % 10 = 5 := by sorry
Prediction:
theorem test_problem : (Finset.sum (Finset.Icc 1 9) (fun i => i^2)) % 10 = 5 := by sorry
These two theorems produce exactly the same numerical result. However, the natural language description specifies computing the squares of the first nine positive integers, while the Label computes the squares of the first ten positive integers. Although this does not change the final value, the semantics are different. Nevertheless, BEq reports that the two theorems are equivalent, which does not match the actual situation.
EPLA-miniF2F #639
NL: How many integers between 15 and 85 are divisible by 20? Show that it is 4
Label:
theorem mathd_numbertheory_12 : Finset.card (Finset.filter (fun x => 20 | x) (Finset.Icc 15 85)) = 4 := by sorry
Prediction:
theorem test_problem : (Finset.filter (fun n => 20 | n) (Finset.range 85 \ Finset.range 16)).card = 4 := by sorry
Although the Label and Prediction produce the same numerical result, the Label is not consistent with the natural language: it additionally includes 15 and 85 in the computation. While this does not affect the numerical result, it is not semantically equivalent to the natural language. On the other hand, the Prediction is strictly equivalent to the natural language. Therefore, the Label and Prediction are not equivalent.
EPLA-miniF2F #641
NL: What is the sum of the units digits of all the multiples of $3$ between 0 and $50$? Show that it is 78
Label:
theorem mathd_numbertheory_447 : ($\mathrm{\Sigma}$ k in Finset.filter (fun x => 3 | x) (Finset.Icc 1 49), k % 10) = 78 := by sorry
Prediction:
theorem test_problem : ((Finset.range 17).sum (fun i => (3 * i) % 10)) = 78 := by sorry
Although the computations for Label and Prediction are the same, Prediction additionally includes 0 and 50, whereas the natural language statement refers to the numbers between 0 and 50, excluding 0 and 50. Therefore, Prediction does not match the natural language. However, BEq will consider Label and Prediction as equivalent.
F.4.3 Header Mixing
EPLA-ProofNet #280
NL: Show that $\sin(\pi/12)$ is an algebraic number.
Label:
open Real
open scoped BigOperators
theorem exercise_12_12 : IsAlgebraic $\mathbb{Q}$ (sin (Real.pi/12)) := by sorry
Prediction:
theorem my_favorite_theorem : IsAlgebraic $\mathbb{Q}$ (Real.sin ($\mathrm{\mathnormal{\pi}}$ / 12)) := by sorry
@@ F.4.3 Header Mixing
In the Prediction, $\pi$ is not defined, so the two theorems are not equivalent. However, if the headers are merged, then $\pi$ becomes defined and the two theorems become equivalent. Since BEq first merges the headers and then performs mutual provability checking, BEq concludes that the two theorems are equivalent, which does not match the actual situation.
F.5 Prompt Templates
This section presents the prompts employed during the autoformalization phase, utilizing Gemini-2.5-Pro (Comanici et al., 2025) and Qwen3-Max (Yang et al., 2025). Subsequently, we detail the prompts for the majority voting phase, which leverages InternLM2-Math-Plus-7B (Ying et al., 2024b) for back-translation and DeepSeek-V3.2-Exp (DeepSeek-AI, 2025) for the NLI consistency check.
Prompt Template for Autoformalization Please autoformalize the following problem in Lean 4 with a header. Use the following theorem names: my_favorite_theorem. {informal_statement} Your code should start with ‘‘‘Lean4 import Mathlib ‘‘‘ You should only output the theorem statement in Lean 4 format, ending with by sorry. You should NOT output the proof.
Prompt Template for Back-Translation [UNUSED_TOKEN_146]user Convert the formal statement into natural language: ‘‘‘lean {formal_statement} ‘‘‘[UNUSED_TOKEN_145]
[UNUSED_TOKEN_146]assistant